Grade 12Physical Chemistry

A buffer of ph 9.26 is made by dissolving x moles of ammonium sulphate and 0.1 mole of ammonia into 100ml solution . If pk(b) of ammonia is 4.74 , then calculate value of x.

Kumar Setu
7 Years agoGrade 12

2 Answers

Sahani Kumar
7 Years ago
pH=
Sahani Kumar
7 Years ago
pH=pKb+log[salt]/[base]      
given: pH=9.26, pKb = 4.74
           [salt]=x mole/L
           [base]=(0.1*1000)/100= 1mol/L
Now 9.26=4.74+log[x]/[1]
         4.52=log[x]
          x=10^-4.52=3.02*10^-5^